# Renewal theory for Brownian motion with stochastically gated targets

### v1

@@ 4 Higher-dimensional renewal equation
$\int_{\partial\Omega}v_{n}^{*}(\mathbf{z},s,\alpha)v_{m}(\mathbf{z},s,\alpha)d\mathbf{z}=\delta_{m,n}.$ (4.18)
We can now solve equation (4.15) by introducing an eigenfunction expansion of $u^{(\epsilon)}$,
$u^{(\epsilon)}(\mathbf{z})=\sum_{m=0}^{\infty}u^{(\epsilon)}_{m}(s,\alpha)v_{m}(\mathbf{z},s,\alpha).$ (4.19)
(The solution picks up a dependence on the switching rate $\alpha$ through the operator ${\mathbb{L}}_{s,\alpha}$.) Substituting equation (4.19) into (4.15) and taking the inner product with the adjoint eigenfunction $v_{n}^{*}(\mathbf{x},s,\alpha)$ yields
$u^{(\epsilon)}_{m}(s,\alpha)=\frac{1}{\epsilon}\frac{{\mathcal{H}}^{(\epsilon)}_{n}(s,\alpha)}{\mu_{n}(s,\alpha)},\quad{\mathcal{H}}^{(\epsilon)}_{n}(s,\alpha)=\int_{\partial\Omega}v_{n}^{*}(\mathbf{z},s,\alpha)h^{(\epsilon)}(\mathbf{z})d\mathbf{z}.$ (4.20)
Combining with equations (4.10), we have the result (assuming we can reverse the order of summation and integrations)
$\widetilde{\rho}^{(\epsilon)}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{y}))=\frac{1}{\epsilon}\int_{\partial\Omega}\bigg[\sum_{n=0}^{\infty}\frac{v_{n}(\mathbf{y},s,\alpha)v_{n}^{*}(\mathbf{z},s,\alpha)}{\mu_{n}(s,\alpha)}\bigg]\widetilde{p}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z}))d\mathbf{z}.$ (4.21)
We can now substitute (4.21) into the renewal equation (4.5) and safely take the limit $\epsilon\rightarrow 0$ since $\widetilde{p}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z}))=O(\epsilon)$. If we also sum equation (4.5) over $k$ and integrate with respect to $\mathbf{x}\in\Omega$, then we obtain the corresponding renewal equation for the Laplace transformed survival probability
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s):=\int_{\Omega}\sum_{k=\pm 1}\widetilde{\rho}_{k|k_{0}}(\mathbf{x},s|\mathbf{x}_{0})d\mathbf{x},$ (4.22)
@@
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s)$ $=\widetilde{Q}(\mathbf{x}_{0},s)+\int_{\partial\Omega}d\mathbf{y}\int_{\partial\Omega}d\mathbf{z}\bigg[\sum_{n=0}^{\infty}\frac{v_{n}(\mathbf{y},s,\alpha)v_{n}^{*}(\mathbf{z},s,\alpha)}{\mu_{n}(s,\alpha)}\bigg]$ (4.23)
$\hskip 85.35826pt\times\bigg[\lim_{\epsilon\rightarrow 0}\frac{\widetilde{Q}(\mathbf{a}_{\epsilon}(\mathbf{z}),s)}{\epsilon}\bigg]\widetilde{J}_{k_{0}}(\mathbf{y},s|\mathbf{x}_{0}).$
@@
$\widetilde{Q}(\mathbf{x}_{0},s)$ $:=\int_{\Omega}\widetilde{p}(\mathbf{x},s|\mathbf{x}_{0})d\mathbf{x}=\frac{1}{s}\bigg[1-\int_{\partial\Omega}\widetilde{J}(\mathbf{y}^{\prime},s|\mathbf{x}_{0})d\mathbf{y}^{\prime}\bigg].$ (4.24a)
@@
$\widetilde{Q}(\mathbf{z}-\epsilon\mathbf{n}(\mathbf{z}),s)$
$=\frac{1}{s}\bigg[1-\int_{\partial\Omega}\widetilde{J}(\mathbf{y}^{\prime},s|\mathbf{z})d\mathbf{y}^{\prime}+\epsilon\int_{\partial\Omega}{\bm{\nabla}}_{\mathbf{z}}J(\mathbf{y}^{\prime},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{z})d\mathbf{y}^{\prime}+O(\epsilon^{2})\bigg]$
$=-\frac{\epsilon D}{s}{\bm{\nabla}}_{\mathbf{z}}\bigg(\int_{\partial\Omega}{\bm{\nabla}}_{\mathbf{y}^{\prime}}\widetilde{G}(\mathbf{y}^{\prime},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{y}^{\prime})d\mathbf{y}^{\prime}\bigg)\cdot\mathbf{n}(\mathbf{z})+O(\epsilon^{2}).$ (4.25)
Substituting into equation (4.23), we have the following double-integral:
${\mathcal{I}}$ $=-\frac{D}{s}\int_{\partial\Omega}d\mathbf{z}\int_{\partial\Omega}d\mathbf{y}^{\prime}v_{n}^{*}(\mathbf{z},s,\alpha){\bm{\nabla}}_{\mathbf{z}}\bigg({\bm{\nabla}}_{\mathbf{y}^{\prime}}\widetilde{G}(\mathbf{y}^{\prime},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{y}^{\prime})\bigg)\cdot\mathbf{n}(\mathbf{z})$
$=\frac{1}{s}\int_{\partial\Omega}d\mathbf{y}^{\prime}\,{\mathbb{L}}^{*}_{s,0}v_{n}^{*}(\mathbf{y}^{\prime},s,\alpha).$ (4.26)
Hence, equation (4.23) can be rewritten as
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s)$ $=\widetilde{Q}(\mathbf{x}_{0},s)+\frac{1}{s}\sum_{n=0}^{\infty}\frac{1}{\mu_{n}(s,\alpha)}\bigg[\int_{\partial\Omega}d\mathbf{y}\,v_{n}(\mathbf{y},s,\alpha)\widetilde{J}_{k_{0}}(\mathbf{y},s|\mathbf{x}_{0})\bigg]$ (4.27)
$\quad\times\bigg[\int_{\partial\Omega}d\mathbf{y}^{\prime}\,{\mathbb{L}}^{*}_{s,0}v_{n}^{*}(\mathbf{y}^{\prime},s,\alpha)\bigg].$
@@ 4.1 Finite interval revisited
Let us briefly return to the example of a finite interval $\Omega=(0,L)$ for which $\partial{\mathcal{U}}=\{0,L\}$. The corresponding Dirichlet Green’s function is given by equation (3.33) and the D-to-N operator is a $2\times 2$ matrix ${\bf L}(s,\alpha)$ with elements
$L_{ab}(s,\alpha)=-\frac{D}{2}(-1)^{a}(-1)^{b}\left.\partial_{x}\partial_{x^{\prime}}[\widetilde{G}(x^{\prime},s|x)+\widetilde{G}(x^{\prime},s+2\alpha|x)\right|_{x^{\prime}=aL,x=bL},$ (4.28)
for $a,b=0,1$. (The factors of 1-1 ensure that the fluxes are directed towards the exterior of the interval.) We thus find that $L_{ab}(s,\alpha)=-(-1)^{a}[{\bf m}(s)+{\bf m}(s+2\alpha)]_{ab}$, see equation (3.39). Moreover, in equation (4.23) we have the simplifications
@@
We thus recover equation (3.31).
@@ 4.2 Stochastically gated sphere
$v_{nm}(\theta,\phi)=\overline{v}_{nm}(\theta,\phi)=\frac{1}{R}Y_{n}^{m}(\theta,\phi),\quad n\geq 0,\ |m|\leq n,$ (4.29)
@@
$Y_{n}^{m}(\theta,\phi)=\sqrt{\frac{2n+1}{4\pi}\frac{(n-m)!}{(n+m)!}}P_{n}^{m}(\cos\theta){\mathrm{e}}^{im\phi},$ (4.30)
@@
$v^{*}_{nm}(\theta,\phi)=\overline{v}_{nm}^{*}(\theta,\phi)=(-1)^{m}\frac{1}{R}Y_{n}^{-m}(\theta,\phi).$ (4.31)
@@
$\mu_{n}(s,\alpha)=\frac{1}{2}\bigg[\sqrt{\frac{s}{D}}\frac{i_{n}^{\prime}(\sqrt{s/D}R)}{i_{n}(\sqrt{s/D}R)}+\sqrt{\frac{s+2\alpha}{D}}\frac{i_{n}^{\prime}(\sqrt{[s+2\alpha]/D}R)}{i_{n}(\sqrt{[s+2\alpha]/D}R)}\bigg],$ (4.32)
@@
$i_{n}(z)=\sqrt{\frac{\pi}{2z}}I_{n+1/2}(z).$ (4.33)
is the modified spherical Bessel function of the first kind. Since the $n$th eigenvalue is independent of $m$, it has a multiplicity $2n+1$. Hence, rewriting equation (4.27) in spherical polar coordinates with $\mathbf{x}_{0}=(r_{0},\theta_{0}.\phi_{0})$ and $r<R$, we obtain the decomposition
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s)=\widetilde{Q}(\mathbf{x}_{0},s)+\frac{R^{2}}{s}\int{\mathcal{D}}(\theta,\phi)\int{\mathcal{D}}(\theta^{\prime},\phi^{\prime})\sum_{n=0}^{\infty}\sum_{m=0}^{n}\frac{\mu_{n}(s,0)}{\mu_{n}(s,\alpha)}(-1)^{m}$ (4.34)
$\hskip 85.35826pt\times Y_{n}^{m}(\theta,\phi)Y_{n}^{-m}(\theta^{\prime},\phi^{\prime})\widetilde{J}_{k_{0}}(R,\theta,\phi,s|\mathbf{x}_{0}),$
@@
$\int{\mathcal{D}}(\theta,\phi):=\int_{0}^{\pi}\sin\theta d\theta\int_{-\pi}^{\pi}d\phi.$ (4.35)
@@
$\int{\mathcal{D}}(\theta^{\prime},\phi^{\prime})Y_{n}^{-m}(\theta^{\prime},\phi^{\prime})=\delta_{n,0}\delta_{m,0}\sqrt{4\pi},$ (4.36)
@@
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s)$ $=\widetilde{Q}(\mathbf{x}_{0},s)+\frac{R^{2}}{s}\frac{\mu_{0}(s,0)}{\mu_{0}(s,\alpha)}\int{\mathcal{D}}(\theta,\phi)\widetilde{J}_{k_{0}}(R,\theta,\phi,s|\mathbf{x}_{0}).$ (4.37)
Finally, noting from equation (4.7) that
$\widetilde{J}_{k_{0}}(R,\theta,\phi,s|\mathbf{x}_{0})$ $=\widetilde{J}(R,\theta,\phi,s|\mathbf{x}_{0})-k_{0}\widetilde{J}(R,\theta,\phi,s+2\alpha|\mathbf{x}_{0})$
$=-\left.D\partial_{r}\bigg(\widetilde{G}(r,\theta,\phi,s|\mathbf{x}_{0})-k_{0}\widetilde{G}(r,\theta,\phi,s+2\alpha|\mathbf{x}_{0})\bigg)\right|_{r=R},$ (4.38)
@@
$-DR^{2}\int{\mathcal{D}}(\theta,\phi)Y_{n}^{m}(\theta,\phi)\partial_{r}\widetilde{G}(R,\theta,\phi,s|\mathbf{x}_{0})=Y_{n}^{m}(\theta_{0},\phi_{0})\frac{i_{n}(\sqrt{s/D}r_{0})}{i_{n}(\sqrt{s/D}R)}$ (4.39)
@@
$\int{\mathcal{D}}(\theta,\phi)\widetilde{J}(R,\theta,\phi,s|\mathbf{x}_{0})=\frac{1}{R^{2}}\frac{i_{0}(\sqrt{s/D}r_{0})}{i_{0}(\sqrt{s/D}R)},\quad i_{0}(z)=\frac{\sinh z}{z}.$ (4.40)
@@
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s)$ $=\widetilde{Q}(\mathbf{x}_{0},s)+\frac{1}{2s}\frac{\mu_{0}(s,0)}{\mu_{0}(s,\alpha)}\bigg[\frac{i_{0}(\sqrt{s/D}r_{0})}{i_{0}(\sqrt{s/D}R)}-k_{0}\frac{i_{0}(\sqrt{[s+2\alpha]D}r_{0})}{i_{0}(\sqrt{[s+2\alpha]/D}R)}\bigg],$ (4.41)
@@
$\widetilde{Q}(\mathbf{x}_{0},s)$ $=\frac{1}{s}\bigg[1-R^{2}\int{\mathcal{D}}(\theta^{\prime},\phi^{\prime})\widetilde{J}(R,\theta^{\prime},\phi^{\prime},s|\mathbf{x}_{0})\bigg]=\frac{1}{s}\bigg[1-\frac{i_{0}(\sqrt{s/D}r_{0})}{i_{0}(\sqrt{s/D}R)}\bigg].$ (4.42)
We now take the limit $s\rightarrow 0$ in equation (4.41) using the small-$z$ expansion $i_{0}(z)=1+z^{2}/6+O(z^{4})$. This yields
${\mathcal{T}}_{k_{0}}(\mathbf{x}_{0})=\frac{R^{2}-r_{0}^{2}}{6D}+\frac{R}{3D}\bigg[\sqrt{\frac{D}{2\alpha}}\frac{i_{0}(\sqrt{2\alpha/D}R)}{i_{0}^{\prime}(\sqrt{2\alpha/D}R)}\bigg]\bigg[1-k_{0}\frac{i_{0}(\sqrt{2\alpha/D}r_{0})}{i_{0}(\sqrt{2\alpha/D}R)}\bigg].$ (4.43)
@@
$\lim_{\alpha\rightarrow 0}{\mathcal{T}}_{-1}(x_{0})=\infty,\quad\lim_{\alpha\rightarrow 0}{\mathcal{T}}_{1}(x_{0})=2\frac{R^{2}-r_{0}^{2}}{6D}.$ (4.44)

### v2

@@ 4 Higher-dimensional renewal equation
$\int_{\partial\Omega}v_{n}^{*}(\mathbf{z},s,\alpha)v_{m}(\mathbf{z},s,\alpha)d\mathbf{z}=\delta_{m,n}.$ (2.4gsajafgouafapatavr)
We can now solve equation (2.4gsajafgouafapatavo) by introducing an eigenfunction expansion of $u^{(\epsilon)}$,
$u^{(\epsilon)}(\mathbf{z})=\sum_{m=0}^{\infty}u^{(\epsilon)}_{m}(s,\alpha)v_{m}(\mathbf{z},s,\alpha).$ (2.4gsajafgouafapatavs)
(The solution picks up a dependence on the switching rate $\alpha$ through the operator ${\mathbb{L}}_{s,\alpha}$.) Substituting equation (2.4gsajafgouafapatavs) into (2.4gsajafgouafapatavo) and taking the inner product with the adjoint eigenfunction $v_{n}^{*}(\mathbf{x},s,\alpha)$ yields
$u^{(\epsilon)}_{m}(s,\alpha)=\frac{1}{\epsilon}\frac{{\mathcal{H}}^{(\epsilon)}_{n}(s,\alpha)}{\mu_{n}(s,\alpha)},\quad{\mathcal{H}}^{(\epsilon)}_{n}(s,\alpha)=\int_{\partial\Omega}v_{n}^{*}(\mathbf{z},s,\alpha)h^{(\epsilon)}(\mathbf{z})d\mathbf{z}.$ (2.4gsajafgouafapatavt)
Combining with equations (2.4gsajafgouafapatavj), we have the result (assuming we can reverse the order of summation and integrations)
$\widetilde{\rho}^{(\epsilon)}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{y}))=\frac{1}{\epsilon}\int_{\partial\Omega}\bigg[\sum_{n=0}^{\infty}\frac{v_{n}(\mathbf{y},s,\alpha)v_{n}^{*}(\mathbf{z},s,\alpha)}{\mu_{n}(s,\alpha)}\bigg]\widetilde{p}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z}))d\mathbf{z}.$ (2.4gsajafgouafapatavu)
We can now substitute (2.4gsajafgouafapatavu) into the renewal equation (2.4gsajafgouafapatave) and safely take the limit $\epsilon\rightarrow 0$ since $\widetilde{p}_{k|-1}(\mathbf{x},s|\mathbf{a}_{\epsilon}(\mathbf{z}))=O(\epsilon)$. If we also sum equation (2.4gsajafgouafapatave) over $k$ and integrate with respect to $\mathbf{x}\in\Omega$, then we obtain the corresponding renewal equation for the Laplace transformed survival probability
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s):=\int_{\Omega}\sum_{k=\pm 1}\widetilde{\rho}_{k|k_{0}}(\mathbf{x},s|\mathbf{x}_{0})d\mathbf{x},$ (2.4gsajafgouafapatavv)
@@
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s)$ (2.4gsajafgouafapatavw)
$=\widetilde{Q}(\mathbf{x}_{0},s)\ +\int_{\partial\Omega}d\mathbf{y}\int_{\partial\Omega}d\mathbf{z}\bigg[\sum_{n=0}^{\infty}\frac{v_{n}(\mathbf{y},s,\alpha)v_{n}^{*}(\mathbf{z},s,\alpha)}{\mu_{n}(s,\alpha)}\bigg]\bigg[\lim_{\epsilon\rightarrow 0}\frac{\widetilde{Q}(\mathbf{a}_{\epsilon}(\mathbf{z}),s)}{\epsilon}\bigg]\widetilde{J}_{k_{0}}(\mathbf{y},s|\mathbf{x}_{0}).$
@@
$\widetilde{Q}(\mathbf{x}_{0},s)$ $:=\int_{\Omega}\widetilde{p}(\mathbf{x},s|\mathbf{x}_{0})d\mathbf{x}=\frac{1}{s}\bigg[1-\int_{\partial\Omega}\widetilde{J}(\mathbf{y}^{\prime},s|\mathbf{x}_{0})d\mathbf{y}^{\prime}\bigg].$ (2.4gsajafgouafapatavxa)
@@
$\widetilde{Q}(\mathbf{z}-\epsilon\mathbf{n}(\mathbf{z}),s)$ $=\frac{1}{s}\bigg[1-\int_{\partial\Omega}\widetilde{J}(\mathbf{y}^{\prime},s|\mathbf{z})d\mathbf{y}^{\prime}+\epsilon\int_{\partial\Omega}{\bm{\nabla}}_{\mathbf{z}}J(\mathbf{y}^{\prime},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{z})d\mathbf{y}^{\prime}+O(\epsilon^{2})\bigg]$ (2.4gsajafgouafapatavxy)
$=-\frac{\epsilon D}{s}{\bm{\nabla}}_{\mathbf{z}}\bigg(\int_{\partial\Omega}{\bm{\nabla}}_{\mathbf{y}^{\prime}}\widetilde{G}(\mathbf{y}^{\prime},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{y}^{\prime})d\mathbf{y}^{\prime}\bigg)\cdot\mathbf{n}(\mathbf{z})+O(\epsilon^{2}).$
Substituting into equation (2.4gsajafgouafapatavw), we have the following double-integral:
${\mathcal{I}}$ $=-\frac{D}{s}\int_{\partial\Omega}d\mathbf{z}\int_{\partial\Omega}d\mathbf{y}^{\prime}v_{n}^{*}(\mathbf{z},s,\alpha){\bm{\nabla}}_{\mathbf{z}}\bigg({\bm{\nabla}}_{\mathbf{y}^{\prime}}\widetilde{G}(\mathbf{y}^{\prime},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{y}^{\prime})\bigg)\cdot\mathbf{n}(\mathbf{z})$ (2.4gsajafgouafapatavxz)
$=\frac{1}{s}\int_{\partial\Omega}d\mathbf{y}^{\prime}\,{\mathbb{L}}^{*}_{s,0}v_{n}^{*}(\mathbf{y}^{\prime},s,\alpha).$
Hence, equation (2.4gsajafgouafapatavw) can be rewritten as
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s)$ $=\widetilde{Q}(\mathbf{x}_{0},s)+\frac{1}{s}\sum_{n=0}^{\infty}\frac{1}{\mu_{n}(s,\alpha)}\bigg[\int_{\partial\Omega}d\mathbf{y}\,v_{n}(\mathbf{y},s,\alpha)\widetilde{J}_{k_{0}}(\mathbf{y},s|\mathbf{x}_{0})\bigg]$
$\quad\hskip 142.26378pt\times\bigg[\int_{\partial\Omega}d\mathbf{y}^{\prime}\,{\mathbb{L}}^{*}_{s,0}v_{n}^{*}(\mathbf{y}^{\prime},s,\alpha)\bigg].$
@@ 4.1 Finite interval revisited
Let us briefly return to the example of a finite interval $\Omega=(0,L)$ for which $\partial{\mathcal{U}}=\{0,L\}$. The corresponding Dirichlet Green’s function is given by equation (2.4gsajafgouafag) and the D-to-N operator is a $2\times 2$ matrix ${\bf L}(s,\alpha)$ with elements
$L_{ab}(s,\alpha)=-\frac{D}{2}(-1)^{a}(-1)^{b}\left.\partial_{x}\partial_{x^{\prime}}[\widetilde{G}(x^{\prime},s|x)+\widetilde{G}(x^{\prime},s+2\alpha|x)\right|_{x^{\prime}=aL,x=bL},$ (2.4gsajafgouafapatavxab)
for $a,b=0,1$. (The factors of 1-1 ensure that the fluxes are directed towards the exterior of the interval.) We thus find that $L_{ab}(s,\alpha)=-(-1)^{a}[{\bf m}(s)+{\bf m}(s+2\alpha)]_{ab}$, see equation (2.4gsajafgouac). Moreover, in equation (2.4gsajafgouafapatavw) we have the simplifications
@@
We thus recover equation (LABEL:ren2gateQ).
@@ 4.2 Stochastically gated sphere
$v_{nm}(\theta,\phi)=\overline{v}_{nm}(\theta,\phi)=\frac{1}{R}Y_{n}^{m}(\theta,\phi),\quad n\geq 0,\ |m|\leq n,$ (2.4gsajafgouafapatavxac)
@@
$Y_{n}^{m}(\theta,\phi)=\sqrt{\frac{2n+1}{4\pi}\frac{(n-m)!}{(n+m)!}}P_{n}^{m}(\cos\theta){\mathrm{e}}^{im\phi},$ (2.4gsajafgouafapatavxad)
@@
$v^{*}_{nm}(\theta,\phi)=\overline{v}_{nm}^{*}(\theta,\phi)=(-1)^{m}\frac{1}{R}Y_{n}^{-m}(\theta,\phi).$ (2.4gsajafgouafapatavxae)
@@
$\mu_{n}(s,\alpha)=\frac{1}{2}\bigg[\sqrt{\frac{s}{D}}\frac{i_{n}^{\prime}(\sqrt{s/D}R)}{i_{n}(\sqrt{s/D}R)}+\sqrt{\frac{s+2\alpha}{D}}\frac{i_{n}^{\prime}(\sqrt{[s+2\alpha]/D}R)}{i_{n}(\sqrt{[s+2\alpha]/D}R)}\bigg],$ (2.4gsajafgouafapatavxaf)
@@
$i_{n}(z)=\sqrt{\frac{\pi}{2z}}I_{n+1/2}(z).$ (2.4gsajafgouafapatavxag)
is the modified spherical Bessel function of the first kind. Since the $n$th eigenvalue is independent of $m$, it has a multiplicity $2n+1$. Hence, rewriting equation (4) in spherical polar coordinates with $\mathbf{x}_{0}=(r_{0},\theta_{0}.\phi_{0})$ and $r<R$, we obtain the decomposition
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s)$ $=\widetilde{Q}(\mathbf{x}_{0},s)+\frac{R^{2}}{s}\int{\mathcal{D}}(\theta,\phi)\int{\mathcal{D}}(\theta^{\prime},\phi^{\prime})\sum_{n=0}^{\infty}\sum_{m=0}^{n}\frac{\mu_{n}(s,0)}{\mu_{n}(s,\alpha)}$
$\qquad\hskip 85.35826pt\times(-1)^{m}Y_{n}^{m}(\theta,\phi)Y_{n}^{-m}(\theta^{\prime},\phi^{\prime})\widetilde{J}_{k_{0}}(R,\theta,\phi,s|\mathbf{x}_{0}),$
@@
$\int{\mathcal{D}}(\theta,\phi):=\int_{0}^{\pi}\sin\theta d\theta\int_{-\pi}^{\pi}d\phi.$ (2.4gsajafgouafapatavxai)
@@
$\int{\mathcal{D}}(\theta^{\prime},\phi^{\prime})Y_{n}^{-m}(\theta^{\prime},\phi^{\prime})=\delta_{n,0}\delta_{m,0}\sqrt{4\pi},$ (2.4gsajafgouafapatavxaj)
@@
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s)$ $=\widetilde{Q}(\mathbf{x}_{0},s)+\frac{R^{2}}{s}\frac{\mu_{0}(s,0)}{\mu_{0}(s,\alpha)}\int{\mathcal{D}}(\theta,\phi)\widetilde{J}_{k_{0}}(R,\theta,\phi,s|\mathbf{x}_{0}).$ (2.4gsajafgouafapatavxak)
Finally, noting from equation (2.4gsajafgouafapatavg) that
$\widetilde{J}_{k_{0}}(R,\theta,\phi,s|\mathbf{x}_{0})$ $=\widetilde{J}(R,\theta,\phi,s|\mathbf{x}_{0})-k_{0}\widetilde{J}(R,\theta,\phi,s+2\alpha|\mathbf{x}_{0})$ (2.4gsajafgouafapatavxal)
$=-\left.D\partial_{r}\bigg(\widetilde{G}(r,\theta,\phi,s|\mathbf{x}_{0})-k_{0}\widetilde{G}(r,\theta,\phi,s+2\alpha|\mathbf{x}_{0})\bigg)\right|_{r=R},$
@@
$-DR^{2}\int{\mathcal{D}}(\theta,\phi)Y_{n}^{m}(\theta,\phi)\partial_{r}\widetilde{G}(R,\theta,\phi,s|\mathbf{x}_{0})=Y_{n}^{m}(\theta_{0},\phi_{0})\frac{i_{n}(\sqrt{s/D}r_{0})}{i_{n}(\sqrt{s/D}R)}$ (2.4gsajafgouafapatavxam)
@@
$\int{\mathcal{D}}(\theta,\phi)\widetilde{J}(R,\theta,\phi,s|\mathbf{x}_{0})=\frac{1}{R^{2}}\frac{i_{0}(\sqrt{s/D}r_{0})}{i_{0}(\sqrt{s/D}R)},\quad i_{0}(z)=\frac{\sinh z}{z}.$ (2.4gsajafgouafapatavxan)
@@
${\mathfrak{Q}}_{k_{0}}(\mathbf{x}_{0},s)$ $=\widetilde{Q}(\mathbf{x}_{0},s)+\frac{1}{2s}\frac{\mu_{0}(s,0)}{\mu_{0}(s,\alpha)}\bigg[\frac{i_{0}(\sqrt{s/D}r_{0})}{i_{0}(\sqrt{s/D}R)}-k_{0}\frac{i_{0}(\sqrt{[s+2\alpha]D}r_{0})}{i_{0}(\sqrt{[s+2\alpha]/D}R)}\bigg],$ (2.4gsajafgouafapatavxao)
@@
$\widetilde{Q}(\mathbf{x}_{0},s)$ $=\frac{1}{s}\bigg[1-R^{2}\int{\mathcal{D}}(\theta^{\prime},\phi^{\prime})\widetilde{J}(R,\theta^{\prime},\phi^{\prime},s|\mathbf{x}_{0})\bigg]=\frac{1}{s}\bigg[1-\frac{i_{0}(\sqrt{s/D}r_{0})}{i_{0}(\sqrt{s/D}R)}\bigg].$ (2.4gsajafgouafapatavxap)
We now take the limit $s\rightarrow 0$ in equation (2.4gsajafgouafapatavxao) using the small-$z$ expansion $i_{0}(z)=1+z^{2}/6+O(z^{4})$. This yields
${\mathcal{T}}_{k_{0}}(\mathbf{x}_{0})=\frac{R^{2}-r_{0}^{2}}{6D}+\frac{R}{3D}\bigg[\sqrt{\frac{D}{2\alpha}}\frac{i_{0}(\sqrt{2\alpha/D}R)}{i_{0}^{\prime}(\sqrt{2\alpha/D}R)}\bigg]\bigg[1-k_{0}\frac{i_{0}(\sqrt{2\alpha/D}r_{0})}{i_{0}(\sqrt{2\alpha/D}R)}\bigg].$ (2.4gsajafgouafapatavxaq)
@@
$\lim_{\alpha\rightarrow 0}{\mathcal{T}}_{-1}(x_{0})=\infty,\quad\lim_{\alpha\rightarrow 0}{\mathcal{T}}_{1}(x_{0})=2\frac{R^{2}-r_{0}^{2}}{6D}.$ (2.4gsajafgouafapatavxar)